OpenStudy (anonymous):

can you help me with this problem 5(x+3)=3x-10

OpenStudy (anonymous):

5x + 15 = 3x - 10 5x - 3x = - 10 - 15 2x = - 25 x = -25/2

OpenStudy (anonymous):

We use DISTRIBUTIVE LAW in the first step and then it is just simplifying

OpenStudy (anonymous):

thanks a lot that helps

OpenStudy (anonymous):

Click medal button if u wish to giv me one!!